40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1463     Entry time: Thu Apr 9 12:23:49 2009
Author: pete 
Type: Update 
Category: Locking 
Subject: tuning ETM common mode 

Pete, Yoichi

Last night, we put the IFO in FP Michelson configuration.  We took transfer functions of CARM and DARM, first using CM excitations directly on the ETMs, and then using modulations of the laser frequency via MC excitation.  We found that there was basically no coupling into DARM using the MC excitation, but that there was coherence in DARM using the ETM excitation.  Therefore, I tuned the ETM common mode in the output matrix.  I did this by taking transfer functions of PD1_Q with PD2_I (see attached plot).  I changed the  drdown_bang script to set C1:LSC-BTMTRX_14 0.98 and C1:LSC-BTMTRX_24 1.02.

Attachment 1: FPMI-DARM-CARM-ETM-fineScan.pdf  4 kB  | Hide | Hide all
FPMI-DARM-CARM-ETM-fineScan.pdf
ELOG V3.1.3-